Please please please help me.
1) A missile is fired into the air from a 100 ft hilltop such that its height above ground after ‘t’ seconds is given by the equation: h(t) = 100 + 800t – 16tsquared
a) Find the time interval for which the missile is at least 9700 ft above ground level.
b) Determine after how many seconds the missile will hit the ground.

Answer by stanbon(48545)

1) A missile is fired into the air from a 100 ft hilltop such that its height above ground after ‘t’ seconds is given by the equation: h(t) = 100 + 800t – 16tsquared
a) Find the time interval for which the missile is at least 9700 ft above ground level.
-16t^2+800t+100=9700
-16t^2+800t-9600 = 0
t^2-50t+600 = 0
(t-20)(t-30)=0
t=30 ; t=30
The missile is at or above 9700 ft between 20 and 30 seconds
after launch.
------------------
b) Determine after how many seconds the missile will hit the ground.
-16t^2+800t+100=0
-4t^2+200t+25 = 0
t = (-200 +- sqrt(200^2-4*-4*25)]/-8
Positive solution:
t = (-200 - sqrt(40400)]/-8
t = 50.12 seconds
========================
Cheers,
Stan H.